OpenStudy (anonymous):

Rational expressions 15xy/x^2-y^2 - X-Y/X+Y

OpenStudy (anonymous):

@sammixboo @whpalmer4 @Nnesha @peachpi

Nnesha (nnesha):

\[\huge\rm \frac{ 15xy }{ x^2-y^2} - \frac{ x-y }{ x+y }\] like this ?

Nnesha (nnesha):

if yes remember difference of square \[\huge\rm a^2 -b^2 = (a+b)(a-b)\] 2nd step ) find common denominator

Nnesha (nnesha):

3rd) cancel some stuff 4th) done!

OpenStudy (anonymous):

@Nnesha Can you show a step by step i've been stuck for hours

OpenStudy (anonymous):

@Nnesha Can you show a step by step i've been stuck for hours

Nnesha (nnesha):

oky i have 5 mints let's see :-) so x ^2 - y^2 = ???

Nnesha (nnesha):

difference of square^

OpenStudy (anonymous):

(x+y)(x-y)?

Nnesha (nnesha):

yes right!

Nnesha (nnesha):

now find common factor

Nnesha (nnesha):

\[\huge\rm \frac{ 15xy - (x-y)(x+y)}{ (x+y)(x-y) }\] now solve x+y are same at both denominator so just write on time and x-y is a common denominator multiply top part with( x+y) denominator of first fraction is same as common denominator so t 15xy stay same

Nnesha (nnesha):

got it ??

Nnesha (nnesha):

now just solve the numerator